# Chengxuan Qian1,  Huaxiu Yao4, Zhengzhong Tu [^1] footnotemark: 1

### v1

@@ 3 Methods
3.1 Preference Generation
Generating high-quality preference data, which includes both accurate ground-truth responses and controlled hallucinated examples, is crucial for effective preference optimization in pre-trained VLMs. Existing methods construct preference data by perturbing ground-truth responses (Zhou et al., 2024a), corrupting visual inputs/embeddings (Deng et al., 2024; Amirloo et al., 2024) to create rejected responses, or refining responses to obtain chosen responses (Chen et al., 2024; Yu et al., 2023a). Refinement produces high quality preference data but comes at a high cost, whereas direct corruption is more scalable yet tends to generate unrealistic hallucinations and fails to produce plausible, natural ones in a controlled manner. To address these limitations, we introduce a novel image retrieval-based pipeline for preference data construction as shown in Figure 3, which consists of three key stages:
3.2 Preference Optimization
w. POVID 88.21 87.16 85.06 10.5495 18.2081 41.5385 40.9302 46.6785
w. SIMA 88.10 87.10 85.03 10.9890 17.6301 43.0549 40.2326 45.2728
w. STIC 89.03 88.20 86.56 12.9670 17.3410 47.2527 34.1860 46.3242
LLaVA-v1.5-7B 66.02 58.18 31.6 50.03 64.1 1510.28 357.85 64.60 3.375
w. POVID 65.98 58.18 31.8 49.80 67.3 1495.91 356.07 64.34 3.625
w. CSR (3Iter) 65.46 57.86 31.6 47.02 68.3 1525.44 365.35 64.08 3.625
w. SIMA 65.83 58.48 32.0 50.04 66.9 1510.33 371.78 64.60 2.5
w. Re-Align 68.10 58.55 32.1 50.06 67.7 1511.79 367.50 64.69 1.375
w. STIC 76.42 63.50 47.3 54.21 81.0 1504.91 308.21 69.16 2.625
w. Re-Align 76.47 64.08 48.3 57.27 81.8 1512.09 318.93 69.42 1.25
4 Experiments
We conduct three categories of experiments to empirically validate the effectiveness of our proposed method. First, we evaluate the ability of Re-Align to mitigate hallucinations and improve generalizability across diverse VQA tasks, demonstrating its consistent superiority over baseline approaches and achieving state-of-the-art performance. Next, we examine Re-Align’s effectiveness in aligning VLMs across various model sizes and architectures, including both text-to-image and unified models, where it delivers substantial performance over vanilla models and existing baselines. Finally, we assess the impact of our proposed rDPO objective in preference optimization, showing that it consistently surpasses standard DPO in aligning VLMs and achieving superior results in both hallucination mitigation and general tasks.
4.1 Re-Align for VLMs Alignment
Datasets
Beslines
We compare our method with several widely adopted alignment frameworks for VLMs, including POVID Zhou et al. (2024a), CSR (Zhou et al., 2024b), SIMA (Wang et al., 2024c), STIC (Deng et al., 2024). For more details on these baselines, please refer to the Appendix.
Methods POPEr POPEp POPEa
Janus-Pro-1B 85.46 85.03 84.13
w. Re-Align 87.53↑2.07 87.33↑2.30 85.86↑1.73
Janus-Pro-7B 88.41 87.30 85.70
w. Re-Align 89.73↑1.32 88.37↑1.07 86.27↑0.57
LLaVA-v1.5-7B 88.14 87.23 85.10
w. POVID 88.21↑0.07 87.16↓0.07 85.06↓0.04
w. CSR (3Iter) 87.83↓0.31 87.00↓0.23 85.00↓0.10
w. SIMA 88.10↓0.04 87.10↓0.13 85.03↓0.07
w. Re-Align 88.65↑0.51 87.43↑0.20 85.16↑0.06
LLaVA-v1.5-13B 88.07 87.53 85.60
w. CSR (3Iter) 88.38↑0.31 87.90↑0.37 85.46↓0.14
w. SIMA 88.04↓0.03 87.40↓0.13 85.40↓0.20
w. Re-Align 90.03↑1.96 89.20↑1.30 86.20↑0.74
LLaVA-v1.6-     Vicuna-7B 88.52 87.63 86.36
w. Re-Align 88.94↑0.42 88.03↑0.40 86.63↑0.27
LLaVA-v1.6-     Vicuna-13B 88.24 87.70 86.43
w. Re-Align 88.79↑0.55 88.10↑0.40 86.60↑0.17
Table 3: Impact of Re-Align across various model scales on POPE.
Methods SQA TextVQA MM-Vet MMEP MMEC MMBench POPEr POPEp POPEa
LLaVA-v1.5-7B 66.02 58.18 31.6 1510.28 357.85 64.60 88.14 87.23 85.10
w. Re-Align (DPO) 66.26 58.24 30.9 1506.49 357.85 64.52 88.18 87.30 85.23
w. Re-Align (rDPO) 68.10 58.55 32.1 1511.79 367.50 64.69 88.65 87.43 85.16
LLaVA-v1.6-Mistral-7B 76.02 63.80 47.6 1494.22 323.92 69.33 88.83 87.93 86.43
w. Re-Align (DPO) 76.07 63.88 46.8 1505.85 316.79 69.24 88.93 88.03 86.47
w. Re-Align (rDPO) 76.47 64.08 48.3 1512.09 318.93 69.42 90.55 89.20 87.03
Table 4: Impact of rDPO across general and hallucination benchmarks for VLMs, and comparisons with baselines.
Experimental Setup
We sample 11k images from the LLaVA-Instruct-150K dataset (Liu et al., 2024a) to construct preference data, as illustrated in Figure 3. These images are initially used to generate QA piars based on image captions and simple VQA tasks using GPT-4o mini (OpenAI, 2024). Furthermore, the images are encoded using clip-vit-large-patch14 (Radford et al., 2021a) to construct the knowledge base for image retrieval. For rejected responses, we use GPT-4o mini to mask the chosen response and all-mpnet-base-v2 (Reimers and Gurevych, 2019) to compute the similarity between the completed masked response and the original chosen response. We use LLaVA-v1.5-7B (Liu et al., 2024a) and LLaVA-v1.6-Mistral-7B (Li et al., 2024b) as our backbone models and perform Re-Align finetuning for $1$ epoch. All evaluations are conducted with a temperature setting of 0, and baseline results are reproduced using the publicly available model weights.
Results
Table 1 shows the performance of Re-Align compared to baseline methods on hallucination benchmarks. Notably, Re-Align achieves the best among the evaluated methods on both POPE and HallusionBench for LLaVA-v1.5-7B (Liu et al., 2024a) and LLaVA-v1.6-Mistral-7B (Li et al., 2024b), highlighting the effectiveness of our approach in mitigating hallucinations of VLMs. As demonstrated in Table 2, Re-Align can provide generally on-par or better performance than the vanilla models and baseline alignment methods on each evaluated general VQA task, ultimately achieving the best overall results. This finding indicates that Re-Align can enhance hallucination mitigation without compromising general performance.
4.2 Scalability and Generalizability
Experimental Setup
  • Unified Models: encoder-decoder architecture that decouples visual encoding for multimodal understanding and generation. In this section, we evaluate Re-Align on Janus-Pro-1B (Chen et al., 2025) and Janus-Pro-7B (Chen et al., 2025).
Results
4.3 Effects of rDPO
Dataset
Due to budget constraints and the need for reproducibility, we have excluded benchmarks that require evaluation by GPT-4 (Achiam et al., 2023). Instead, we focus on the following tasks: ScienceQA (Lu et al., 2022), TextVQA (Singh et al., 2019), MM-Vet (Yu et al., 2023b), VisWiz (Gurari et al., 2018), LLaVABench (Liu, 2023), MME (Fu et al., 2023), MMBench (Liu et al., 2024c), and POPE (Li et al., 2023c).
Experimental Setup
The experimental setup follows the same setting as VLMs alignment experiments, with the exception of the direct optimization objectives. To further explore the impact of our proposed rDPO, we conduct experiments on the same constructed preference dataset using the standard DPO (Rafailov et al., 2024) during the one-epoch finetuning process.
Results
5 Discussions
Discussion with mDPO
$\mathcal{L}_{\text{CoDPO}}=-\mathbb{E}_{(x,v,y_{w},y_{l})\sim\mathcal{D}}$
In contrast, Re-Align extends and enhances this approach by incorporating a more semantically meaningful visual preference pair. Instead of relying solely on random crops, Re-Align retrieves a relevant image from the same dataset that corresponds to the original input. This retrieval-based augmentation provides a stronger contrastive signal, improving the model’s ability to discern fine-grained visual details and reducing spurious correlations. Moreover, beyond mitigating hallucinations in VLMs, Re-Align has been demonstrated that it also significantly enhances performance on general VQA tasks.
Segment-level Preference
6 Related Work
Reinforcement Learning from Human Feedback
Vision Language Models
Large Vision Language Models (VLMs) (Li et al., 2022, 2023a; Liu et al., 2024a; Li et al., 2024b; Meta, 2024; Bai et al., 2023; Wang et al., 2024b; Lu et al., 2024; Wu et al., 2024) extended the understanding and reasoning capabilities of Large Language Models (LLMs) (Devlin et al., 2018; Radford et al., 2019; Brown et al., 2020; Team et al., 2023; Roziere et al., 2023; Touvron et al., 2023a, b; Raffel et al., 2020; Yang et al., 2024; Team, 2024) into the visual domain. By integrating vision encoders, such as CLIP (Radford et al., 2021b), image patches are first converted into embeddings and then projected to align with text embedding space, unlocking unprecedented cross-modal applications in the real world, such as biomedical imaging (Moor et al., 2023; Li et al., 2024a), autonomous systems (Shao et al., 2024; Tian et al., 2024; Sima et al., 2023; Xing et al., 2024b), and robotics (Rana et al., 2023; Kim et al., 2024).
Alignment of Vision Language Models
Current VLMs often suffer from hallucinations, producing inaccurate or misleading information that fails to accurately represent the content of the provided image (Zhu et al., 2024; Bai et al., 2024). Such misalignments can have catastrophic consequences when these models are deployed in real-world scenarios (Xing et al., 2024a). To address cross-modality hallucinations, recent research has primarily focused on applying direct preference optimization (Deng et al., 2024; Zhou et al., 2024a; Fang et al., 2024; Zhou et al., 2024b; Guo et al., 2024; Chen et al., 2024; Wang et al., 2024c; Yu et al., 2024b; Li et al., 2023b; Wang et al., 2024a) or contrastive learning (Sarkar et al., 2024) on the curated datasets with preference signals, and utilizing model editing techniques (Liu et al., 2024b; Yu et al., 2024a).
7 Conclusion
In this paper, a novel framework, Re-Align, for aligning VLMs to mitigate hallucinations is proposed. Our approach leverages image retrieval to deliberately induce segment-level hallucinations, thereby generating plausible and natural preference signals in a controlled manner. By integrating the retrieved images, a dual-preference dataset that encompasses both textual and visual cues is curated. Furthermore, we propose the rDPO objective, an extension of DPO that includes an additional visual preference optimization objective, to enhance the alignment process with valuable visual preference signals. Comprehensive empirical results from a range of general VQA and hallucination benchmarks demonstrate that Re-Align effectively reduces hallucinations in VLMs while enhancing their overall performance. Moreover, it demonstrates superior scalability across various model architectures and sizes.
Limitations
Although Re-Align has demonstrated superior performance on both hallucination and general VQA benchmarks, it does not always achieve state-of-the-art results on general tasks; in some cases, its performance is even worse than that of vanilla VLMs. Future research could explore strategies to eliminate this alignment tax or or identify an optimal balance for this trade-off.
The potential risks of this work align with the general challenges of RLHF alignment. As more powerful alignment techniques are developed, they may inadvertently empower adversarial approaches that exploit these models, potentially leading to unfair or discriminatory outputs. Meanwhile, these adversarial strategies can be used to generate negative samples, which can ultimately contribute to the development of more robust and reliable VLMs over time.
Appendix A Details of the Evaluated Baselines
We compare our proposed method with following alignment frameworks for VLMs:
@@ Appendix A Details of the Evaluated Baselines
Appendix B Prompts used for Preference Data Construction
@@
Appendix C Examples of Preference Pair
Table 5 and 6 provide examples of the constructed preference data for the VQA and image captioning, and each data sample contains textual instruction, input image, retrieved image, chosen response, and rejected response.
@@ Appendix C Examples of Preference Pair
Appendix D Response Examples
@@
Appendix E Licenses
@@
Appendix F Experimental Cost
The cost for curation the preference dataset by using GPT-4o mini (OpenAI, 2024) cost approximately \$90 in total.The evaluation of HallusionBench and LLaVABench using GPT-4 (Achiam et al., 2023) incurred an approximate total cost of \$30.
Appendix G Computational Cost
@@
Appendix H Hyperparameter Setting
@@
Appendix I Social Impacts

### v2

@@ Appendix A Overview of Re-Align
Appendix B Details of the Evaluated Baselines
We compare our proposed method with the following alignment frameworks for VLMs:
  • LLaVA-RLHF (Sun et al., 2023): conducts SFT on for updating the projector only and then PPO on the preference data collected from human annotators.
@@ Appendix B Details of the Evaluated Baselines
  • mDPO (Wang et al., 2024a): finetunes the model with conditional preference optimization, which incorporates an additional objective to account for image-level preferences and a reward anchor that forces the reward to be positive for chosen responses.
Appendix C Prompts used for Preference Data Construction
@@
Appendix D Examples of Preference Pair
Table 4 and 5 provide examples of the constructed preference data for the VQA and image captioning, and each data sample contains textual instruction, input image, retrieved image, chosen response, and rejected response.
@@ Appendix D Examples of Preference Pair
Appendix E Response Examples
@@
Appendix F Additional Discussions
Performance Variations on General VQA tasks
While Re-Align consistently delivers the best performance on hallucination benchmarks across all backbone models, it may not achieve the top result for every general VQA benchmark. The variations in performance on general VQA tasks are primarily due to the alignment tax, a well-known phenomenon in RLHF, where alignment can sometimes lead to a decline in the model’s ability to retain pretraining knowledge. Notably, this trade-off is not unique to Re-Align; as shown in Table 2, several baselines even underperform compared to the vanilla VLMs on general VQA tasks.
Segment-level Preference
Appendix G Licenses
@@
Appendix H Experimental Cost
The cost for curating the preference dataset by using GPT-4o mini (OpenAI, 2024) cost approximately \$90 in total.The evaluation of HallusionBench and LLaVABench using GPT-4 (Achiam et al., 2023) incurred an approximate total cost of \$30.
Appendix I Computational Cost
@@
Appendix J Hyperparameter Setting
@@
Appendix K Social Impacts